OpenStudy (anonymous):

if volume= 314 and h=3 find the value of r helpppp

OpenStudy (anonymous):

\[V=\frac{r^2 \times \pi \times h}{3}\]\[314=\frac{3r^2 \pi}{3}\]\[314\times3=3r^2 \pi\]\[942=3r^2 \pi\]\[ r^2=\frac{942}{3\pi}\]\[r^2=99.9\]\[r=\sqrt{100} \]\[r=10\]

OpenStudy (anonymous):

Brit_95 I can't find prerivos problem I miss square, I need correct it

OpenStudy (anonymous):

http://openstudy.com/groups/Mathematics?invite=sudhanshu_kmr#/groups/mathematics/updates/4ddd9c0dee2c8b0bf28b41e8 check new answer, I just corect it I miss square